OpenStudy (anonymous):

Laura has moved to a new apartment. Her schoolbooks comprising of different subjects are mixed in a bag during the move. Four books are of mathematics, three are English, and six are science. If Laura opens the bag and selects books at random, find the given probability. P(1 science and 2 mathematics books)

OpenStudy (anonymous):

it is not explicitly mentioned, but I assume she is taking three books at random. now, probability is positive outcomes/all outcomes all outcomes; out of 13 books (4+3+6) she picks 3; so that is 13c3 positive outcomes; we know she has to draw 1 science and 2 mathematics in order for draw to be successful; she has to draw 2 math books out of 4 and 1 science book out of 6, that means positive outcome odd is (4c2)*(6c1) \[P = \frac{ \left(\begin{matrix}4 \\ 2\end{matrix}\right)*\left(\begin{matrix}6 \\ 1\end{matrix}\right) }{ \left(\begin{matrix}13 \\ 3\end{matrix}\right) }\]

OpenStudy (anonymous):

Thank you soo much for you're help.

OpenStudy (anonymous):

np;)